OpenStudy (anonymous):

Please help! Find the following infinite sum, if it exists. (LOOK AT RESPONSE) Thank you!

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty}2\left( -1/2 \right)^{n-1}\]

OpenStudy (anonymous):

O.k. So how would I find the infinite sum?

OpenStudy (anonymous):

i forgot sum upto infinity for geometric progression :(

OpenStudy (anonymous):

i thinks its a/(1-r) where a is first term, r is the common ration.. so here a = 2/2= 1 and r = 1/2

OpenStudy (anonymous):

sorry here a= 2

OpenStudy (anonymous):

So it does not exist?

OpenStudy (anonymous):

NO , it does not exist

OpenStudy (anonymous):

Thank you imranmeah91!